OpenStudy (iwanttogotostanford):

HELP PLEASE< AM I RIGHT HERE? :

OpenStudy (iwanttogotostanford):

i think that I don't know what to do last, i know the first steps but not the lasts

OpenStudy (iwanttogotostanford):

i think it is even but again not sure

OpenStudy (iwanttogotostanford):

@volleyballlover55 @Sadlalkdja12111312 @mathmate

OpenStudy (iwanttogotostanford):

@Nnesha

OpenStudy (iwanttogotostanford):

I know I am supposed to replace x with -x and then see if the function stays the same or not, and with this one I got this ar: f(-x)=-x + 12/-x

OpenStudy (mathmate):

So do you conclude that f(-x) does not equal f(x) or do you think they are equal?

OpenStudy (iwanttogotostanford):

thats the part I am not sure about because couldn't -x cancel the other -x out

OpenStudy (iwanttogotostanford):

in the other side of the function? and then it would be back to how it was originally

OpenStudy (mathmate):

no, the -x in f(-x) cannot be cancelled with the -x on the right hand side. The way to look at is is the following, if you're not sure. f(x)=x+1/x f(-x)=-x-1/x If they are equal, when you subtract one from the other, you should get a zero, right?

OpenStudy (iwanttogotostanford):

yes

OpenStudy (mathmate):

So let's do f(x)-f(-x) =x+1/x - (-x -1/x) =? can you work that out, taking care about the sign, and remember to distribute.

OpenStudy (iwanttogotostanford):

x+1/x^2 im sorr that was really confusing and I'm not sure i did it right

OpenStudy (iwanttogotostanford):

in the end though i worked it out and i think it will be even

OpenStudy (mathmate):

if we continue the calculation, we have x+1/x - (-x) - (-1/x) =x+1/x+x+1/x =2x+2/x which is not a zero. Since we say that f(-x) does not equal f(-x), what can you conclude?

OpenStudy (iwanttogotostanford):

that it is in fact an odd function. @mathmate

OpenStudy (iwanttogotostanford):

since they don't equal....

OpenStudy (mathmate):

Unfortunately, there are functions that are neither odd nor even. So you have to check if it is an odd function before you can say that. As you know, an odd function will satisfy f(x)=-f(-x). Since you said f(-x)=-x-1/x Can you tell me if f(-x) is the same as -f(x)?

OpenStudy (iwanttogotostanford):

no it isn't so it IS neither.. sorry

OpenStudy (iwanttogotostanford):

i forgot it could be neither, thats why i was confused

OpenStudy (mathmate):

Not really, see we can factor out the negative sign. -x-1/x = -(x+1/x) which is the negative of f(x), so therefore f(-x)=-f(x). What kind of function would that be?

OpenStudy (iwanttogotostanford):

even

OpenStudy (iwanttogotostanford):

since they are the same

OpenStudy (mathmate):

Even: f(-x)=f(x) Odd: f(-x)=-f(x). Try again!

OpenStudy (iwanttogotostanford):

odd!

OpenStudy (mathmate):

Yes, because we showed that f(-x) = -x -1/x = -(x+1/x) = -f(x), so it is odd.

OpenStudy (iwanttogotostanford):

i said that first and i think you said it was wrong. Anyways, thank you so much! I understand.

OpenStudy (mathmate):

If I missed it, I'm sorry about it. You're welcome! :)